OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2)and perpendicular to y = -x + 6. A.) y = -x + 6 B.) y = x - 6 C.) y = x + 2 D.) y = -x - 2

OpenStudy (mertsj):

What is the slope of the given line?

OpenStudy (anonymous):

It never gave me a slope line.

OpenStudy (mertsj):

It asks for a line perpendicular to y=-x+6 What is the slope of y = -x+6?

OpenStudy (anonymous):

y = 6 + -1x?

OpenStudy (mertsj):